OpenStudy (anonymous):

Suppose an employer agreed to pay you a penny for your first day’s work, then double your wages every day thereafter for 30 days. Using a calculator, find how much the employer would pay you on your 15th day. (Hint: Use the expression 2n-1 to help model this situation.)

OpenStudy (fibonaccichick666):

Can you write the sum using a riemann sum?

OpenStudy (anonymous):

I don't know...

OpenStudy (fibonaccichick666):

have you heard of a riemann sum?

OpenStudy (anonymous):

no

OpenStudy (fibonaccichick666):

ok, can we use excel and do it the easy way?

OpenStudy (anonymous):

sure

OpenStudy (fibonaccichick666):

so just so you know how to check it, then we can do the math way. Excel, input .01 into cell A1. A2=2*A1 A3=2*A2 then use the autofill handle down to A30

OpenStudy (fibonaccichick666):

oh but select cells A1-A3 before you autofill

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

sorry don't select A1

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

So, that is the excel way to find the answer

OpenStudy (fibonaccichick666):

let's do the math way

OpenStudy (anonymous):

ok

OpenStudy (fibonaccichick666):

if n=1 pay=?

OpenStudy (anonymous):

ok....

OpenStudy (fibonaccichick666):

?= I want you to answer

OpenStudy (anonymous):

i don't know... a penny?

OpenStudy (fibonaccichick666):

it's given and yes. Did you do the excel sheet for reference here?

OpenStudy (anonymous):

No, I don't have it on this computer

OpenStudy (fibonaccichick666):

well you have two options, write out the calculations 15 times or find the limit of the summation. I unfortunately just looked at the time and have to be up in 4 hours, night

OpenStudy (fibonaccichick666):

@e.mccormick Could you help for me?

OpenStudy (anonymous):

$327.68